OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation instandard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

Please Help!

OpenStudy (anonymous):

first find the slope between these two points....

OpenStudy (lalaly):

Do you know how to find the slope?

OpenStudy (anonymous):

Slope is rise over run right?

OpenStudy (anonymous):

oops This is the wrong problem

OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation instandard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

That is what I meant to put

OpenStudy (campbell_st):

 the slope is rise over run m = -2/2 m = 1 then the equation of the line is in point slope form is y = mx + b you can substitute m y = -x + b pick a point, say E and substitute x = 4, y = 3 to find b 3 = -1*4 + b b = 7 the equation is y = -x + 7